OpenStudy (richyw):

having trouble with a series expansion. \[f(z)=\frac{z+1}{z-1}\]I need series expansion in both |z|<1 and 1<|z|<∞

OpenStudy (richyw):

for |z|<1 I rewrite as \[-(z+1)\frac{1}{1-z}=\frac{-z}{1-z}+\frac{1}{1+z}\]

OpenStudy (richyw):

oops missing a negative sign there. anyways I get\[=-\sum^\infty_{n=0}z^{n+1}-\sum^\infty_{n=0}z^n\]

OpenStudy (richyw):

for |z|<1. Now Apparently I need to replace n by n-1 in the first series. Can someone explain to me why I do this?

OpenStudy (anonymous):

There's more than one possible answer, but here's what you could also do for \(|z|<1\): \[\frac{z+1}{z-1}=\frac{z-1+2}{z-1}=1-2\sum_{n=0}^\infty z^n\] What you did: \[\frac{z+1}{z-1}=-\frac{z}{1-z}-\frac{1}{1-z}=-\sum_{n=0}^\infty z^{n+1}-\sum_{n=0}^\infty z^n\] Changing \(n\) to \(n-1\) requires changing the starting index so that you have \[\sum_{n=0}^\infty z^{n}=1+z+z^2+z^3+\cdots\\ \sum_{n=1}^\infty z^{n-1}=1+z+z^2+z^3+\cdots\] As for why? I don't see a good reason for it. I like to consolidate and combine the series whenever I can, but changing the index is counterproductive in that regard.

OpenStudy (richyw):

ah, I missed that my text changed the starting index. thanks.

OpenStudy (richyw):

but are you not missing a negative sign? wouldn't it be \[-1-2\sum_{n=0}^\infty z^n\]

OpenStudy (richyw):

oh wait nvm. you also started at n = 0. ok thanks.